A boy throws a ball straight into the air. It reaches the highest point of its flight in 4 seconds. How fast was the ball going when it left the boy’s hand?


1
Expert's answer
2021-08-19T14:23:26-0400

"v=v_0-gt,"

"v_0=v+gt=gt=39.2~\frac ms."